Please show and explain all work for each questions

What decimal number does the bit pattern 11001100 represent if it is a:
• [1 pts] unsigned integer?
• [1 pts] sign-magnitude integer?
• [1 pts] two's complement integer?

4 - What decimal number does the bit pattern 00110011 represent if it is a:
• [1 pts] unsigned integer?
• [1 pts] sign-magnitude integer?
• [1 pts] two's complement integer?

5 - [2 pts] What is the binary representation of the decimal number 5.5 assuming the IEE 754 single precision format?

6 - What decimal number does the bit pattern 0xC0B00000 represent if it is:
• [2 pts] A two's complement integer?
• [2 pts] An unsigned integer?
• [2 pts] A floating point number assuming the IEE 754 single precision format
Answer #1

3).11001100

a). unsigned integer

In unsigned integer there is no negative number.

So, 11001100 = 204

b). sign-magnitude integer

In sign magnitude form the MSB bit tells the sign of number(negative if 0 or positive if 1) and rest tells the magnitude.

11001100 is negative number as MSB is 1 and it's magnitude is 1001100

So, number in sign magnitude form is -76

c).two's complement form

In this form,if MSB is 1 then number is negative else number is positive.

11001100 is negative as it's MSB is 1.

Magnitude of 11001100 = 1's complement of 11001100 +1 = 00110011 +1 = 00110100 = 52

So,11001100 in 2's complement is -52.

4).00110011

a). Unsigned integer

00110011 = 51

b).sign magnitude integer

00110011 is positive as it's MSB is 0.

Magnitude is decimal of 0110011 = 51

So,00110011 in sign magnitude form is -51.

c).2's complement form

Since,number is positive the magnitude is simply decimal equivalent of binary 00110011.

So,00110011 in 2's complement form is +51.

5).5.5